The perimeter of a rectangle is 54. The length is 3 less than 2 times the widht. What is the lenght?
Answer by ewatrrr(24785)


Hi
Let x and (2x-3) represent the width and length respectively
Question states***
2(2x-3) + 2x = 54units
Solving for x
6x = 60
x =10, the width. The length is 17 (20 -3)
CHECKING our Answer***
2*17 + 2*10 = 34 + 20 = 54units